Grade 10Magical Mathematics[Interesting Approach]

N IS A NATURAL NUMBER,1 is written after the five digit number to make it as asix digit number,which is thrice the same number if one is written after it,find the number

kishore
11 Years agoGrade 10

1 Answer

Shobhit Varshney
11 Years ago
Dear student,

10N + 1 = 3(105 + N)
10N + 1 = 3*105 + 3N
7N = 299999
N = 42857
….….….….….….….….….….….….….….….….….…....
Thanks